OpenStudy (anonymous):

wats a^2+b^2 formula someone plz tell

OpenStudy (vishweshshrimali5):

there is no such formula but u can derive it \[(a+b)^2=a^2+b^2+2ab\] \[(a+b)^2-2ab=a^2+b^2\] or \[(a-b)^2=a^2+b^2-2ab\] \[(a-b)^2+2ab=a^2+b^2\]